JANE IS THREE TIMES AS OLD AS HER SISTER. IN THREE YEARS, SHE WILL BE 2 YEARS MORE THAN TWICE OF HER SISTER'S AGE. WHAT ARE THEIR PRESENT AGE? THANK U VERY MUCH...

Answer by checkley71(8403)

JANE=3X
SISTER=X
(3X+3)=2(X+3)+2
3X+3=2X+6+2
3X-2X=6+2-3
X=5 AGE OF THE SISTER NOW.
3*5=15 JANE'S AGE NOW.PROOF:
(3*5+3)=2(5+3)+2
15+3=2*8+2
18=16+2
18=18